# GelSLAM: A Real-time, High-Fidelity, and Robust 3D Tactile SLAM System

### v1

@@ VI-E Large Object Reconstruction using GelBelt
We qualitatively evaluate the reconstruction of a tree trunk with a diameter of approximately 190 mm by scanning around 130 mm of its length using GelBelt. The sensor was rolled along the trunk surface, moving both vertically and radially, and completed nine full rounds around the trunk. Figure 16 shows the resulting reconstruction. The reconstructed surface mesh closely matches the real texture of the trunk, including visible lines and cracks. We observe slight shrinkage in the cylindrical shape near the two ends of the reconstructed mesh, suggesting that minor global distortion can occur in some cases for GelSLAM.
@@
VII Conclusions and Discussion
@@
In this work, we introduce GelSLAM, a novel tactile-only system for long-horizon object tracking and high-fidelity object-level 3D reconstruction from tactile sequences. By leveraging differential representations, GelSLAM robustly captures spatial relationships between tactile readings, even under challenging conditions such as large rotations, minimal overlap, and low surface texture. These capabilities support loop closure and tactile relocalization, enabling the system to correct drift and maintain global spatial consistency over time. As a result, GelSLAM overcomes the inherently local nature of touch and achieves long-horizon tracking as well as building coherent global reconstructions with detailed geometry. In experiments, we show that GelSLAM runs online and maintains accurate tracking across a wide range of everyday objects, effectively correcting drift over long horizons. We also demonstrate, both qualitatively and quantitatively, that GelSLAM reconstructs detailed object-level 3D geometry with high accuracy. Using GelBelt, we further show that our system can reconstruct large objects such as a tree trunk while preserving fine surface details. To the best of our knowledge, GelSLAM is the first tactile-only system to achieve robust and accurate object-level 3D reconstruction at this level of fidelity.
Despite GelSLAM’s strong performance, several challenges remain and suggest promising directions for future work. One limitation is its difficulty with extremely low-texture objects, where loops are hard to detect. For objects with highly repetitive patterns, the GNC optimization can provide partial support, but may still fail or even break correctly detected loops, suggesting the need for more reliable algorithms in these challenging cases. Another constraint of GelSLAM is the requirement to touch the entire object surface, which is time-consuming. This limitation could be addressed by using sparse tactile inputs, with untouched regions filled in based on prior knowledge or through the integration of vision. In addition, the contact mask detection method we use sometimes requires manual tuning of the height threshold, which could be replaced with more adaptive methods [48].
Our current system supports a flexible, in-the-wild setup where a human can operate it without requiring a robot. This makes it practical for real-world use, but can lead to redundant exploration, as it is difficult for humans to identify which regions remain untouched. A natural extension is to enable autonomous robotic exploration that actively plans where to touch next and when to revisit uncertain areas to improve efficiency. We also believe that GelSLAM’s long-horizon object tracking capability could enhance robotic manipulation performance. GelSLAM opens a new direction in tactile perception by extending touch from a local sensing modality to one capable of global, long-horizon spatial understanding. We believe it has the potential to reshape how robots use touch for manipulation tasks.

### v2

@@ VI-E Large Object Reconstruction using GelBelt
We qualitatively evaluate the reconstruction of a tree trunk with a diameter of approximately 190 mm by scanning around 130 mm of its length using GelBelt. The sensor was rolled along the trunk surface, moving both vertically and radially, and completed nine full rounds around the trunk. Fig. 18 shows the resulting reconstruction. The reconstructed surface mesh closely matches the real texture of the trunk, including visible lines and cracks. We observe slight shrinkage near the two ends of the reconstructed cylindrical mesh. This is likely because the middle regions are constrained by pose graph edges from multiple directions, while the ends are constrained mainly by one-sided pose graph connections, making them more susceptible to global distortion.
@@
VII Discussion
Our experiments demonstrate that GelSLAM achieves real-time, robust long-horizon tracking, together with high-fidelity reconstruction, across a wide range of objects. At the same time, we identify several challenging conditions that motivate future research directions.
VII-A Challenging Conditions and Partial Solutions
For objects with extremely low surface texture, GelSLAM can struggle with loop detection, leading to pose drift, relocalization failure, and disconnected or locally discontinuous regions in the reconstructed mesh. Fig. 19 shows the GelSLAM reconstruction of an egg, a representative near-textureless object. Only the largest connected component of the reconstructed mesh is shown. Although the entire surface is scanned, failure to relocalize after contact loss leads to disconnected mesh segments. In addition, the reconstructed mesh exhibits local discontinuities, which arise from uncorrected pose misalignment when regions are revisited but loop detection is missed. Despite these issues, GelSLAM still achieves a partial reconstruction, with $11$ tracking sessions, $221$ keyframes, and $70$ detected loops forming the largest connected mesh, suggesting meaningful relocalization and loop detection capability despite minimal surface texture. In general, touch alone is unlikely to resolve this limitation for extremely low-textured objects. Incorporating complementary sensing modalities, such as vision, may provide the global context needed for improved loop detection.
Figure 19: Challenging scenario. For extremely low-textured objects, GelSLAM may fail to detect loop closures, resulting in fragmented and locally discontinuous reconstructions.
Another challenge arises for objects with highly repetitive surface patterns, where loop detection may incorrectly associate frames from different locations. Although such repetition is uncommon in GelSight sensing due to its high resolution, it can occur in industrially manufactured objects and lead to outlier loop closures that significantly degrade pose estimation and reconstruction. As an initial approach, we propose to apply a GNC-based pose graph optimization [60] to reject outlier loops, which can improve reconstruction quality in these cases as shown in Fig. 20. However, the increased computational cost and the fact that GNC can occasionally reject valid loops indicate that better solutions are still needed.
Figure 20: Challenging scenario. For objects with repetitive patterns, GelSLAM may suffer from outlier loop closures. Applying GNC-based pose graph optimization filters these outliers and significantly improves reconstruction quality.
VII-B Online Reconstruction
GelSLAM can run in real-time for at least 10 minutes without noticeable performance drop. Since the primary real-time use case of GelSLAM is object pose tracking for manipulation, which typically lasts well under 10 minutes, this time span is more than sufficient in practice. For reconstruction, most applications do not require real-time mesh output and can therefore be performed offline. For the rare cases requiring online reconstruction, Fig. 21 shows that online results closely match offline reconstruction for 7-minute (peanut) and 12-minute (small rock) scans, but introduce artifacts for a 19-minute scan (large rock) due to loop skipping.
Figure 21: Online vs. offline GelSLAM reconstruction. Online reconstruction closely matches offline results for 7- and 12-minute scans, but introduces artifacts for a longer 19-minute scan. In practice, reconstruction rarely needs to be real-time, and offline reconstruction meets most use cases.
VII-C Future Directions
Several extensions could further improve the efficiency, robustness, and accuracy of GelSLAM. One promising direction is to reduce the need to touch the entire object surface by inferring untouched regions using prior knowledge or completing them with visual information. In addition, GelSLAM estimates contact masks via simple height thresholding, as in NormalFlow [12] for its simplicity. One could adopt learning-based contact mask estimation methods [50] for potentially better performance.
Another direction is to incorporate priors or explicit inertial sensing from IMUs to improve pose tracking accuracy. Example priors include assuming approximately planar motion during contact, imposing a constant velocity prior, and incorporating object dynamics models. Incorporating these into the pose graph optimization could further improve the accuracy of GelSLAM.
Finally, while GelSLAM supports flexible, in-the-wild human operation without requiring a robotic platform, manual scanning can lead to redundant exploration, as it is difficult to identify which regions remain untouched. A natural extension is autonomous tactile exploration, where a robot actively plans the next touching locations and revisits uncertain regions to improve scanning efficiency and accuracy. Extending the system to multi-fingered tactile sensing offers another avenue to accelerate coverage and further improve efficiency.
VIII Conclusions
@@
In this work, we introduce GelSLAM, a novel tactile-only system for long-horizon object tracking and high-fidelity object-level 3D reconstruction from tactile sequences. Our main contribution is a set of tactile-specific SLAM components, including NormalFlow failure detection, keyframe selection, and loop detection, together with their system-level integration into a well-engineered SLAM system. By leveraging differential representations, GelSLAM robustly captures spatial relationships between tactile readings, even under challenging conditions such as large rotations, minimal overlap, and low surface texture. These capabilities support robust loop closure and tactile relocalization, enabling the system to correct drift and maintain global spatial consistency over time. Unlike prior systems limited to hundreds of frames and can frequently include false loop detections, GelSLAM scales to tens of thousands of frames with zero false loops. As a result, GelSLAM overcomes the inherently local nature of touch and achieves long-horizon tracking as well as building coherent global reconstructions with detailed geometry.
In experiments, we show that GelSLAM runs online and maintains accurate tracking across a wide range of everyday objects, effectively correcting drift over long horizons. We also demonstrate, both qualitatively and quantitatively, that GelSLAM reconstructs detailed object-level 3D geometry with high accuracy. Using GelBelt, we further show that our system can reconstruct large objects such as a tree trunk while preserving fine surface details. To the best of our knowledge, GelSLAM is the first tactile-only system to achieve robust and accurate object-level 3D reconstruction at this level of fidelity.
To conclude, GelSLAM’s real-time, long-horizon object tracking capability can enhance robot manipulation performance and potentially reshape how robots utilize touch. Beyond tracking, its high-fidelity reconstruction performance demonstrates that tactile sensing can serve as a practical reconstruction modality, particularly for applications requiring high precision under heavy occlusion or involving transparent objects. We believe GelSLAM opens a new direction in tactile perception and manipulation by extending touch from a local sensing modality to one capable of global, long-horizon spatial understanding.